OpenStudy (anonymous):

A wedding cake is prepared using 3 individual cakes. The first is a rectangular prism with a length inches, a width 9 inches, and a height 4 inches. The second is a cylinder with a diameter 8 inches and a height 2 inches. The third is a cube with side length 2 inches. What is the approximate total volume of the cake? Use 3.14 for . Round to the nearest whole number.

OpenStudy (anonymous):

length of prisma ?

OpenStudy (anonymous):

for cube 2^3 = 8 cylinder = 1/4 *3.14*8*8*2 = 100.48 and calculate for the prism ..

OpenStudy (anonymous):

volume prism = length x width x height and adding the volume

OpenStudy (anonymous):

the lenth is 13 sorry

OpenStudy (anonymous):

just do some calculation :)